# Comment on “Galvano-rotational effect induced by electroweak interactions in pulsars”

### v1

@@ III The corrected energy spectrum for Dvornikov’s paper
$\gamma=\gamma_{\sigma}\equiv\left(l+\frac{\sigma}{2}\text{sgn}(V_{L})-\frac{1}{2}\right),\ \ \beta=\beta_{\sigma}\equiv\frac{1}{4|V_{L}|\omega}\left[E^{2}+2E(J_{z}\omega-V_{L})+(V_{L})^{2}+J^{2}_{z}\omega^{2}+V_{L}\omega\sigma-p^{2}_{z}\right],$ (16)
or better:
$\beta\equiv\frac{1}{4|V_{L}|\omega}\left[(E-V_{L}+J_{z}\omega)^{2}+2\omega V_{L}\left(J_{z}+\frac{\sigma}{2}\right)-p^{2}_{z}\right].$ (17)
Now, let us solve Eq. (15) through of the asymptotic behavior/limit for $\rho\to 0$ and $\rho\to\infty$ (i.e. two asymptotic behaviors) [3, 4, 7, 9, 10, 11, 12, 13, 14, 16, 17]. Therefore, for $\rho\to 0$, we have:
$\left[\rho\frac{d^{2}}{d\rho^{2}}+\frac{d}{d\rho}-\frac{\gamma^{2}}{4\rho}\right]\varphi_{\sigma}(\rho)\approx 0,$ (18)
whose solution is given by $\varphi_{\sigma}(\rho)=\rho^{|\gamma|/2}$.
Now, for $\rho\to\infty$, we have:
$\left[\rho\frac{d^{2}}{d\rho^{2}}-\frac{\rho}{4}\right]\varphi_{\sigma}(\rho)\approx 0,$ (19)
whose solution is given by $\varphi_{\sigma}(\rho)=e^{-\rho/2}$.
In this way, a regular solution for Eq. (15) can be written as:
$\varphi_{\sigma}(\rho)=Ce^{-\rho/2}\rho^{|\gamma|/2}F_{\sigma}(\rho),$ (20)
where $C$ is a normalization constant and $F_{\sigma}(\rho)$ is a unknown function to be determined. In particular, to be a physically allowed solution (or a normalizable solution), $\varphi_{\sigma}(\rho)$ must satisfy two boundary conditions of the system, i.e. vanish at the origin and infinity, such as $\varphi_{\sigma}(\rho\to 0)=\varphi_{\sigma}(\rho\to\infty)=0$.
So, replacing the solution (20) in (15), we obtain a second-order differential equation for $F_{\sigma}(\rho)$:
$\left[\rho\frac{d^{2}}{d\rho^{2}}+\left(\left|\gamma\right|+1-\rho\right)\frac{d}{d\rho}-\left(\frac{\left|\gamma\right|+1}{2}-\beta\right)\right]F_{\sigma}(\rho)=0,$ (21)
or yet (“Dvornikov-type form”)
$\left[\rho\frac{d^{2}}{d\rho^{2}}+\left(|\gamma|+1-\rho\right)\frac{d}{d\rho}+\kappa\right]F_{\sigma}(\rho)=0,\ \ \kappa\equiv\left(-\frac{|\gamma|+1}{2}+\beta\right).$ (22)
According to Refs. [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17], the equation above is an associated Laguerre equation (or a confluent hypergeometric equation), whose solution are the associated Laguerre polynomials (or confluent hypergeometric functions). Consequently, the parameter $\kappa$ must be equal to a positive integer, that is, $\kappa=N$, where $N=0,1,2,\ldots$ is the radial quantum number (i.e. a quantum number associated with the radial/spatial part of the spinor). In fact, the correct thing to say would be that the last term of (21) is a negative integer, but since the minus sign has already been incorporated into $\kappa$, it amounts to the same thing/result. Therefore, from the condition $\kappa=N$ (a quantization condition), we obtain the following corrected energy spectrum for Dvornikov’s paper:
$E=V_{L}-J_{z}\omega\pm\mathcal{E},\ \ \mathcal{E}=\sqrt{p^{2}_{z}+4|V_{L}|N_{\sigma}\omega},\ \ (J_{z}=(1/2-l)\text{sgn}(V_{L})),$ (23)
@@
$N_{\sigma}\equiv\left(N+\frac{1}{2}+\frac{|J_{z}-\frac{\sigma}{2}|}{2}-\textbf{sgn}(V_{L})\frac{(J_{z}+\frac{\sigma}{2})}{2}\right)=\left(N+\frac{1-\sigma\text{sgn}(V_{L})}{2}+\frac{|J_{z}-\frac{\sigma}{2}|}{2}-\text{sgn}(V_{L})\frac{(J_{z}-\frac{\sigma}{2})}{2}\right),$ (24)
being $N_{\sigma}=N_{eff}=N_{total}$ a type of “effective or total quantum number” (since it depends on all others), and is similar to that of Refs. [10, 11, 12, 13, 14, 28, 29], and we use $V_{L}=\text{sgn}(V_{L})|V_{L}|$. In Table 1, we have the values of $N_{\sigma}$ for eight possible combinations of $\sigma$, $\text{sgn}(V_{L})$, and $J_{z}$. So, we see clearly/explicitly that for $J_{z}>0$ (neutrino with positive angular momentum), $\sigma=+1$ (upper component of the Dirac spinor) and $V_{L}>0$, or $J_{z}<0$ (neutrino with negative angular momentum), $\sigma=-1$ (lower component of the Dirac spinor) and $V_{L}<0$, we recover the particular energy spectrum of Dvornikov’s paper with $m^{2}=0$ (as it should be) [2].
Table 1: Values of $N_{\sigma}$ for eight possible combinations of $\sigma$, $\text{sgn}(V_{L})$, and $J_{z}$.
$N_{\sigma}$ $\sigma$ $\text{sgn}(V_{L})$ $J_{z}$
$N_{\sigma}=N$ $\sigma=+1$ $\text{sgn}(V_{L})=+1$ $J_{z}>0$
$N_{\sigma}=N+\frac{1}{2}-J_{z}$ $\sigma=+1$ $\text{sgn}(V_{L})=+1$ $J_{z}<0$
$N_{\sigma}=N+\frac{1}{2}+J_{z}$ $\sigma=+1$ $\text{sgn}(V_{L})=-1$ $J_{z}>0$
$N_{\sigma}=N+1$ $\sigma=+1$ $\text{sgn}(V_{L})=-1$ $J_{z}<0$
$N_{\sigma}=N+1$ $\sigma=-1$ $\text{sgn}(V_{L})=+1$ $J_{z}>0$
$N_{\sigma}=N+\frac{1}{2}-J_{z}$ $\sigma=-1$ $\text{sgn}(V_{L})=+1$ $J_{z}<0$
$N_{\sigma}=N+\frac{1}{2}+J_{z}$ $\sigma=-1$ $\text{sgn}(V_{L})=-1$ $J_{z}>0$
$N_{\sigma}=N$ $\sigma=-1$ $\text{sgn}(V_{L})=-1$ $J_{z}<0$
IV Final remarks
In the present comment, we obtain the corrected energy spectrum for Dvornikov’s paper [2], that is, a spectrum in which the matter term, given by $V_{L}$, is “tied” in both the radial and angular quantum numbers, given by $N$ and $J_{z}$ (where $N\geq 0$ and $J_{z}\gtrless 0$), and not only to $N$ as shown by Dvornikov. To be more specific, the matter term in Dvornikov’s paper [2] that we refer to is that found inside a square root and not that found outside, in which it is a free term. In particular, the error committed by Dvornikov arose from an incorrectly solved second-order differential equation (after having made a change of variable), where only one asymptotic behavior was considered, in which the scalar function vanishes at infinity (the correct is to vanish at both origin and infinity). Besides, such differential equation also had another error, i.e. since Dvornikov considered ultrarelativistic particles where $m\to 0$ (“massless neutrinos”), it implies that all terms of this equation containing $m$ should vanish (or be zero); however, he zeroed one term and kept another (which could not have happened). Therefore, using as a base several papers in the literature on the Dirac equation (in curved space-times, with and without rotation), we solve in detail the corrected Dvornikov equation and obtain the corrected spectrum for the Dirac neutrinos. Finally, we verified that for $J_{z}>0$ (positive angular momentum) with $\sigma=+1$ (upper component of the Dirac spinor) and $V_{L}>0$, or $J_{z}<0$ (negative angular momentum) with $\sigma=-1$ (lower component of the Dirac spinor) and $V_{L}<0$, we recover the particular energy spectrum of Dvornikov’s paper with $m^{2}=0$ (as it should be).

### v2

@@ II Quick review of the main steps that Dvornikov took to obtain the energy spectrum
where $\tilde{Q}^{a}=q^{a}-q_{\text{eff}}\tilde{A}^{a}_{\text{eff}}$, being $q^{a}=\left(E+J_{z}\omega-V_{V},-i\partial_{r},0,p_{z}\right)$, $\tilde{V}=V_{A}\gamma^{0}\gamma^{5}$, and $\tilde{A}^{a}_{\text{eff}}=\left(0,\frac{i}{2q_{\text{eff}}r},-\frac{1}{q_{\text{eff}}}\frac{J_{z}}{r},0\right)$.
Now, defining $\psi_{r}\equiv[\gamma^{a}\tilde{Q}_{a}+m-\tilde{V}]\Phi$, we can obtain from Eq. (26) a second-order differential equation for the spinor $\Phi$, given as follows
$\left[\left(\partial_{r}+\frac{1}{2r}\right)^{2}+(E+J_{z}\omega-V_{V})^{2}-\frac{J_{z}^{2}}{r^{2}}-2V_{A}\gamma^{5}(E+J_{z}\omega-V_{V})+\frac{J_{z}}{r^{2}}\Sigma_{3}-p^{2}_{z}+V_{A}^{2}+2mV_{A}\gamma^{0}\gamma^{5}-m^{2}\right]\Phi=0,$ (27)
or better
$\left[\partial_{r}^{2}+\frac{1}{r}\partial_{r}-\frac{(J_{z}-\frac{\Sigma_{3}}{2})^{2}}{r^{2}}+(E+J_{z}\omega-V_{V})^{2}-2V_{A}\gamma^{5}(E+J_{z}\omega-V_{V})-p^{2}_{z}+V_{A}^{2}+2mV_{A}\gamma^{0}\gamma^{5}-m^{2}\right]\Phi=0,$ (28)
where $\Sigma_{3}=\gamma^{0}\gamma^{3}\gamma^{5}$ and $\Sigma_{3}^{2}=1$.
Here, such as stated/considered by Dvornikov [2], the solution of Eq. (28) can be found for ultrarelativistic particles ($m\to 0$); consequently, it implies that all terms of the equation containing $m$ must vanish (or be zero). If not (i.e. $m\neq 0$), then to solve the equation, it must first be diagonalized (which would be something very complex) [38]. In other words, the product $\gamma^{0}\gamma^{5}$ is anti-diagonal (this can be easily shown using (7)), i.e. $\gamma^{0}\gamma^{5}$=anti-diag(+1,1)(+1,-1). Therefore, here, we consider the limit $m\to 0$. Besides, we also consider left-handed neutrinos where $(1+\gamma^{5})\psi=0$ or $V_{R}=0$, which implies $V_{V,A}=V_{L}/2$. In this way, using these considerations and $\Phi$ written as $\Phi=\upsilon\varphi$ (with $\Sigma_{3}\upsilon=\sigma\upsilon$ and $\gamma^{5}\upsilon=+\upsilon$, and $\sigma=\pm 1$), Eq. (28) becomes
$\left[\partial_{r}^{2}+\frac{1}{r}\partial_{r}-\frac{(J_{z}-\frac{\sigma}{2})^{2}}{r^{2}}+\left(E+J_{z}\omega-\frac{V_{L}}{2}\right)^{2}-V_{L}\left(E+J_{z}\omega-\frac{V_{L}}{2}\right)+\frac{V^{2}_{L}}{4}-p^{2}_{z}\right]\varphi_{\sigma}=0,$ (29)
or better
$\left[\frac{d^{2}}{dr}+\frac{1}{r}\frac{d}{dr}-\frac{A_{\sigma}^{2}}{r^{2}}+B^{2}\right]\varphi_{\sigma}(r)=0,$ (30)
@@
$A_{\sigma}\equiv\Big|J_{z}-\frac{\sigma}{2}\Big|\geq 0,\ \ B\equiv\sqrt{(E+J_{z}\omega-V_{L})^{2}-p^{2}_{z}}>0.$ (31)
IV The correct energy spectrum for a neutrino interacting with matter in a rotating frame
According to Refs. [39, 40, 41, 42, 43, 44], Eq. (30) is the well-known Bessel equation (a type of second-order differential equation), whose solution (“radial wave function”) is given by
$\varphi_{\sigma}(r)=c_{1}J_{A_{\sigma}}(Br)+c_{2}Y_{A_{\sigma}}(Br),$ (32)
where $J_{A_{s}}(Br)$ and $Y_{A_{s}}(Br)$ are the Bessel functions of first and second kinds, and $c_{1}$ and $c_{2}$ are arbitrary constants. So, as at the origin $(r=0)$ the Bessel function of first kind is finite/regular while the Bessel function of second kind is infinite/irregular, implies that: $c_{1}\neq 0$ and $c_{2}=0$ [39, 40, 41, 42, 43, 44]. Therefore, the solution (32) becomes
$\varphi_{\sigma}(r)=c_{1}J_{A_{\sigma}}(Br).$ (33)
where must satisfy $\varphi_{\sigma}(r\to 0)=0$, i.e., this is our first boundary condition (null function at the origin).
Now, we would like to restrict the motion of the neutrino in a region where a hard-wall confining potential is present, i.e., a finite distance where the function is also null. So, according to Refs. [39, 40, 41, 42, 43, 44], the following boundary condition (our second boundary condition) describes this kind of confinement
$\varphi_{\sigma}(r=r_{0})=0,$ (34)
which means that the function vanishes at a fixed radius $r_{0}$, with $0<r_{0}<\infty$. According to Refs. [39, 40, 41, 42, 43, 44], a plausible physical motivation/argument for using this second boundary condition (or hard-wall confining potential) is that it allows obtaining from a Bessel equation the quantization of the energy (i.e. to obtain the bound states). In particular, we can say that this boundary condition is similar to the finite/infinite square potential well [45], which confines a non-relativistic particle in a finite region in space and, consequently, generates the energy’s quantization (i.e. the confinement leads to quantization). As we will see later, our spectrum will depend on $1/r^{2}_{0}$ just as it happens in the non-relativistic spectrum for a particle subject to the finite/infinite square potential well (in fact, since here we will have a relativistic spectrum, $1/r^{2}_{0}$ will be within the square root).
Besides, assuming that $Br_{0}\gg 1$, the Bessel function (33) can be written as follows [39, 40, 41, 42, 43, 44]
$J_{A_{\sigma}}(Br_{0})\to\sqrt{\frac{2}{\pi Br_{0}}}\cos\left(Br_{0}-\frac{A_{\sigma}\pi}{2}-\frac{\pi}{4}\right).$ (35)
Consequently, the boundary condition (34) implies that the argument of (35) must satisfy the following relation [39, 40, 41, 42, 43, 44]
$\left(Br_{0}-\frac{A_{\sigma}\pi}{2}-\frac{\pi}{4}\right)=(2N+1)\frac{\pi}{2},\ \ (N=0,1,2,\ldots),$ (36)
where $n$ is a quantum number, sometimes called the radial quantum number (since it arises from a radial differential equation). Therefore, using the quantization condition (36) with (31), we obtain the following energy spectrum for a neutrino interacting with matter in a rotating frame (and subject to a hard-wall confining potential)
$V_{L}+\left(l-\frac{1}{2}\right)\omega\text{sgn}(V_{L})\pm\mathcal{\tilde{E}},\ \ \mathcal{\tilde{E}}=\sqrt{p^{2}_{z}+\frac{\pi^{2}}{r^{2}_{0}}\left(N+\frac{|J_{z}-m_{s}|}{2}+\frac{3}{4}\right)^{2}}$. (37)
[…]
V Final remarks
In the present comment, we show that the fundamental equation worked by Dvornikov [2] in his paper is incorrect, which is given by the Dirac equation for a neutrino interacting with matter in a rotating frame (or curved space-time). In particular, Dvornikov [2] incorrectly wrote/defined the interaction term (effective potentials or background matter) in a curved space-time. That is, such potentials should also depend on the vierbein; however, this was not what happened. Consequently, one of his main results (if not the main one), namely the relativistic energy spectrum (or quantized energy eigenvalues), is also incorrect (actually partially incorrect). So, starting from the correct Dirac equation with effective potentials in a curved space-time, we obtain the correct energy spectrum for a neutrino interacting with matter in a rotating frame. We observe that the square root of this spectrum is very different from that obtained by Dvornikov [2]. In fact, in our case, the spectrum was obtained from a Bessel equation subject to a hard-wall confining potential (analogous to the well-known finite/infinite square potential well), while the Dvornikov spectrum [2] was obtained from an associated Laguerre equation. However, we note that the only correct thing in the Dvornikov spectrum [2] are the terms outside the square root, which agree with those of the correct spectrum.